OpenStudy (anonymous):

~~~~~~~FAN AND MEDAL!~~~~~~~~~~~~~~~~~~~~ What is the value of the y variable in the solution to the following system of equations? 2x + y = 5 x + 3y = 0 −1 1 3 −3

OpenStudy (anonymous):

2x + y = 5 x + 3y = 0 ----> x=-3y 2x + y = 5 substitute 2(-3y) + y = 5 -6y+y=5 -5y=5 y=-1

OpenStudy (anonymous):

So it is A.

OpenStudy (anonymous):

Thank You!!

OpenStudy (anonymous):

Anytime!